OpenStudy (anonymous):

How do I derive computational formula for standard deviation from definition using elementary algebra?

OpenStudy (amistre64):

statistical standard deviation from elementary algebra? HA!!

OpenStudy (saifoo.khan):

\begin{array}l\color{#787299}{\text{A}}\color{#8482a8}{\text{m}}\color{#771444}{\text{i}}\color{#491918}{\text{s}}\color{#454615}{\text{t}}\color{#646534}{\text{r}}\color{#11aa71}{\text{e}}\color{#982713}{\text{!}}\color{#384724}{\text{!}}\color{#814849}{\text{}}\end{array}

OpenStudy (amistre64):

standard deviation in simplest terms is just the average distance stuff is from the mean

OpenStudy (anonymous):

so would it just be square root of sum of means divided by n?

OpenStudy (amistre64):

for a population yes; that is a good assessment

OpenStudy (anonymous):

hmm seems too easy haha

OpenStudy (amistre64):

sum of the distances from the mean for every data point

OpenStudy (saifoo.khan):

http://www.answers.com/topic/standard-deviation

OpenStudy (amistre64):

the distance is squared to get positive values to sum up and then is divided by the number in the population and square rooted to account for the squaring to begin with

OpenStudy (amistre64):

when playing with samples you run into degrees of freedom and other adjustments :)

OpenStudy (anonymous):

I'm confused....how would i write this as an equation other than how standard deviation is usually expressed?

OpenStudy (amistre64):

have you tried using celtic runes? :)

OpenStudy (anonymous):

no what is that

OpenStudy (amistre64):

its chicken scratch that the norsemen used to write in lol

OpenStudy (amistre64):

\[\sigma=\sqrt{\frac{\sum{(x-\bar x)^2}}{N}}\]

OpenStudy (amistre64):

xbar should be mu tho :) \[\bar x = \mu\]

OpenStudy (anonymous):

yaa thats the formula I am familiar with, but I am supposed to find another way to write it using algebra I guess...its an extra credit question

OpenStudy (amistre64):

all algebra is moves things around

OpenStudy (amistre64):

\[\sigma^2(N)=\sum(x-\mu)^2\] maybe?

OpenStudy (anonymous):

hmm its worth try haha

OpenStudy (anonymous):

thanks!

OpenStudy (amistre64):

good luck with it :)

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

I am supose to do the same thing and this does not sound right

OpenStudy (anonymous):

are you in stats summer school ivc??

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

well you're in my class then. ya i don't get what he is looking for..

OpenStudy (anonymous):

As I asked for Frodolokalike same question, what definition were you given in class?

OpenStudy (anonymous):

just the one in the book, but i think computational formula means what you would input into your calculator...

OpenStudy (anonymous):

I think that's what computational formula means as well (or into a computer). I mean its very straightforward if you just have a small population or sample. Was the definition in the book "square root of variance"?

OpenStudy (anonymous):

yaa that or it gave you the equation sum of (x-mean)^2 all divided by 2

OpenStudy (anonymous):

so the computational formula for variance is here http://en.wikipedia.org/wiki/Variance#Computational_formula ....maybe you could just square that

OpenStudy (anonymous):

i think its s= variance ^2

OpenStudy (anonymous):

oh jk its not

OpenStudy (anonymous):

variance is s^2

OpenStudy (anonymous):

the computational formula is the formula in the book is labeled formula 3-5

OpenStudy (anonymous):

oh i see it...so would that be the answer? seems too easy hah

OpenStudy (anonymous):

i dont think it is... he wants us to define the forumula using elementary algebra

OpenStudy (anonymous):

I think you have already done part of it above.... sqrt(variance), so what's variance? Plug that in. Whats expected value? plug that in....

OpenStudy (anonymous):

I'm so lost

OpenStudy (anonymous):

it almost sounds like a poof like heres the answer figure out how to get that answer using algebra

OpenStudy (anonymous):

yaa I have asked a few people already and still can't figure it out

OpenStudy (anonymous):

I think between you, you can come up with an answer to this question. Sleep on it.

OpenStudy (anonymous):

its due tomarrow at 11

OpenStudy (anonymous):

Have a look at the first section, Basic examples, of http://en.wikipedia.org/wiki/Standard_deviation See if that gives you any ideas.

OpenStudy (anonymous):

sarah91 have u had any progress

OpenStudy (anonymous):

not really I if I dont find anything better I'm just going to write the computational formula to find variance and take the square root of it